OpenStudy (anonymous):

Solve the equation Twenty-five minus four b equals negative six b plus forty-seven for b. A)b equals negative eleven B)b equals negative two and two tenths C)b equals two and two tenths D)b equals eleven

OpenStudy (anonymous):

First, write the equation: \[25 - 4b = -6b + 47\]

OpenStudy (anonymous):

Next, rearrange the equation and combine like terms. Add 6b to both sides. Subtract 25 from both sides. This will put all b terms on the left, and your constants on the right. \[25 - 4b + 6b - 25 = -6b + 47 + 6b - 25\] \[2b = 22\] Solve for b. \[b = \frac{22}{2}\] Simplify. \[b = 11\]